# Radial and vertical constraints on the icy origin of HCO in the HD 163296 Protoplanetary Disk

### v1

@@ Abstract
HCO is a small organic molecule widely detected in protoplanetary disks. As a precursor to grain-surface formation of CHOH, HCO is considered an important precursor of O-bearing organic molecules that are locked in ices. Still, since gas-phase reactions can also form HCO, there remains an open question on the channels by which organics form in disks, and how much the grain versus the gas pathways impact the overall organic reservoir. We present spectrally and spatially resolved Atacama Large Millimeter/submillimeter Array (ALMA) observations of several ortho- and para-HCO transitions toward the bright protoplanetary disk around the Herbig Ae star HD 163296. We derive column density, excitation temperature, and ortho-to-para ratio (OPR) radial profiles for HCO, as well as disk-averaged values of cm, K, and , respectively. We empirically determine the vertical structure of the emission, finding vertical heights of . From the profiles, we find a relatively constant with radius, but still consistent with among the uncertainties, a secondary increase of in the outer disk, and low values that decrease with disk radius. Our resulting radial, vertical, and OPR constraints suggest an increased UV penetration beyond the dust millimeter edge, consistent with an icy origin but also with cold gas-phase chemistry. This Herbig disk contrasts previous results for the T Tauri disk, TW Hya, which had a larger contribution from cold gas-phase chemistry. More observations of other sources are needed to disentangle the dominant formation pathway of HCO in protoplanetary disks.
@@ 1 Introduction
Although complex hydrocarbons and N-bearing COMs are commonly detected in protoplanetary disks, O-bearing COMs detections are scarce (see Öberg et al., 2023, and references therein) because most of them are expected to be locked in the ice mantles on the surface of dust grains (Öberg et al., 2011; Boogert et al., 2015). Even methanol (CHOH), one of the simplest O-bearing COMs and considered a building block in forming more complex species (Öberg et al., 2009), has had relatively few detections in disks. Except for TW Hya (Walsh et al., 2016), the closest protoplanetary disk, CHOH detections are typically restricted to peculiar sources with specific warm conditions (e.g., van ’t Hoff et al., 2018; Lee et al., 2019; Podio et al., 2020) or cavities directly exposed to the radiation of the host star (e.g., van der Marel et al., 2021; Booth et al., 2021, 2023), consistent with an icy origin. Thus, the need arises for an abundant gas-phase molecule capable of tracing the O-rich organic content in protoplanetary disks.
@@
According to previous work (Walsh et al., 2014; Öberg et al., 2017), the small organic molecule formaldehyde (HCO) has the potential to be a tracer of the frozen content in dust grains, since it can efficiently form on ices by the successive hydrogenation of CO, which also leads to the formation of CHOH and other COMs (e.g., Watanabe & Kouchi, 2002; Chuang et al., 2017) and is more accessible to detect in the gas-phase due to lower desorption energies compared to CHOH (Carney et al., 2019). Still, HCO can also be formed through gas-phase reactions (e.g., Fockenberg & Preses, 2002; Atkinson et al., 2006), and thus, it cannot be used as a direct tracer of dust grain chemical products if there are no constraints on the contribution from each formation pathway (Loomis et al., 2015; Terwisscha van Scheltinga et al., 2021).
@@
Consequently, HCO has been extensively studied in multiple protoplanetary disks, mostly around low-mass T Tauri stars. As different regions of the disk are expected to have distinct types of chemistry (Öberg & Bergin, 2021), most studies have focused on resolving the radial distribution of HCO (e.g., Loomis et al., 2015; Öberg et al., 2017; Kastner et al., 2018; Podio et al., 2019; Garufi et al., 2020; Guzmán et al., 2021; Booth et al., 2023). Nevertheless, the disk vertical structure also has an important role in HCO chemistry, which can be constrained by observing edge-on sources (e.g., Podio et al., 2020), using geometrical methods (e.g., Paneque-Carreño et al., 2023), or determining the excitation conditions by observing multiple HCO lines (e.g., Pegues et al., 2020).
@@
Parameter Value Reference Distance pc Gaia Collaboration et al. (2018) InclinationaaThe geometric parameters of the disk are derived from the dust millimeter continuum. deg Huang et al. (2018) P.A.aaThe geometric parameters of the disk are derived from the dust millimeter continuum. deg Huang et al. (2018) Teague et al. (2021) Fairlamb et al. (2015) Stellar type A1 Fairlamb et al. (2015) km s Teague et al. (2019)
@@
Moreover, HCO has two isomeric forms depending on the hydrogen nuclei spin alignment: ortho-HCO (parallel alignment) and para-HCO (anti-parallel alignment). The ratio of the two isomers, known as the ortho-to-para ratio (OPR), has also been considered a possible way to discriminate between HCO formation pathways (e.g., Guzmán et al., 2011). However, the interpretation of derived OPRs is still a matter of discussion (Guzmán et al., 2018; Terwisscha van Scheltinga et al., 2021, see Section 4.2.3 for further details). Currently, the excitation conditions and the OPR of HCO have been well-resolved only for the T Tauri disk TW Hya (Terwisscha van Scheltinga et al., 2021), thanks to Atacama Large Millimeter/submillimeter Array (ALMA) high angular resolution observations of several ortho and para lines of HCO. Therefore, further multiple-line, spatially-resolved studies of p-HCO and o-HCO are needed to gain a representative understanding of HCO chemistry in protoplanetary disks.
@@
Due to its large size ( au, Law et al., 2021a), proximity, and high luminosity (see Table 1), the Herbig Ae disk HD 163296 is an ideal source for a detailed study of HCO chemistry. Currently, only two p-HCO transitions have been spatially resolved in HD 163296 (Carney et al., 2017; Pegues et al., 2020; Guzmán et al., 2021). Still, additional bright p-HCO and o-HCO lines can be potentially resolved (Guzmán et al., 2018), thus covering a broad range of excitation energies to obtain well-constrained estimates of the excitation conditions and the OPR of HCO. The disk also contains a well-defined CO snowline (Qi et al., 2015) and dust millimeter edge (Law et al., 2021a), which makes HD 163296 a perfect target to see how HCO chemistry works in more massive Herbig Ae/Be sources and compare it to the chemistry in the colder T Tauri, TW Hya disk.
@@
In this work, we analyze the emission of multiple o-HCO and p-HCO transitions in HD 163296 resolved at angular resolution with ALMA. We determine empirically the radial and vertical distribution of HCO, as well as its excitation conditions and OPR as a function of radius. The observations and data reduction are described in Section 2. The results on the spatial distribution, excitation conditions and OPR are presented in Section 3, while their implications in the formation of HCO in protoplanetary disks are discussed in Section 4. The conclusions are summarized in Section 5.
@@
HCO Line Baselines Chan. Width BeamaaThe robust parameter used for Briggs weighting and a UV taper, both applied in the CLEAN process, were computed independently for each transition to obtain a common circularized beam of for all lines (see Appendix A). Beam PA Chan. rmsbbThe channel rms is calculated using the central of the pixels in the first and last channels of the data cubes. Mom. Zero rmsccThe moment zero rms is defined by the bootstrapping procedure described in Section 3.2. Int. Flux DensityddThe integrated flux density is measured within an outer radius of and integrating the emission between to km s. The uncertainties do not include the calibration uncertainty. (m) (km s) () () (mJy beam) (mJy beam km s) (mJy km s) ALMA Band 3 o-HCO 11footnotemark: ALMA Band 6 p-HCO 11footnotemark: o-HCO 22footnotemark: ALMA Band 7 p-HCO 22footnotemark: p-HCO 22footnotemark: p-HCO 22footnotemark: o-HCO 22footnotemark:
@@ 2.1 Observational Details
We present ALMA observations of several HCO lines in Band 3, 6, and 7 in the protoplanetary disk HD 163296111HD 163296 protoplanetary disk coordinates are ; .. A detailed description of the Band 6 and 7 observations (project 2016.1.00884.S, PI: V.V. Guzmán) is given in Carney et al. (2019). In summary, the Band 6 observations were carried out with the 12 m array in two different configurations between 2016 November and 2017 March, whereas in the case of Band 7 the short baselines executions were observed with the Atacama Compact Array (ACA) on 2016 October and the long baselines executions were done with the 12m array on 2017 April. Two different spectral setups were used: one targeting the o-HCO rotational transition centered at GHz (Band 6), and other targeting several p-HCO and o-HCO rotational transitions (, , , and ) centered at GHz (Band 7). Transitions of other molecules were also covered within the above setups, such as CHOH (Carney et al., 2019) and CN (Bergner et al., 2021).
Additionally, these observations were complemented with publicly available data from the MAPS ALMA Large Program (project 2018.1.01055.L, PI: K.I. Öberg). All the observational details of these data can be found in Öberg et al. (2021). In particular, we used the detected p-HCO rotational transition at GHz (Band 6) presented by Guzmán et al. (2021), and the non-detected o-HCO rotational line at GHz (Band 3). All the transitions used in this work are summarized in Table 2 along with their respective project codes.
@@ 2.2 Data Reduction
The data reduction and posterior imaging were performed using the Common Astronomy Software Application (CASA222https://casa.nrao.edu/) version 6.5.2.26 (McMullin et al., 2007; CASA Team et al., 2022). The data from 2016.1.00884.S were initially processed through the standard pipeline provided by the Joint ALMA Observatory (JAO), and then self-calibrated to improve the signal-to-noise ratio following a methodology similar to that used in the DSHARP ALMA Large Program (see Andrews et al., 2018, for more details). Essentially, we created pseudo-continuum visibilities by flagging the line emission in the observed spectral windows, and then we aligned the different measurement sets (short and long baselines separately) into a common phase center by using the CASA functions fixvis and fixplanets. Afterward, the self-calibration procedure can be summarized as follows: we first applied the phase-only self-calibration to the short baseline visibilities, followed by a phase-amplitude self-calibration. Finally, we repeated the same procedures for the combined measurement sets (short + long baselines).
@@
The calibrated visibilities were imaged through the CLEAN algorithm (Högbom, 1974) implemented by the tclean task. We used a Briggs weighting and multiscale deconvolver (Cornwell, 2008) with scales of 0, 5, 15, and 25 pixels, where the pixel size corresponds to approximately 1/7 of the beam size. The CLEAN mask was defined using the keplerian_mask package (Teague, 2020) to account for the Keplerian rotation of the disk, using the same stellar and disk parameters of HD 163296 as in MAPS (see Öberg et al., 2021, values are summarized in Table 1), ensuring that we covered all disk emission at different velocities. Regarding continuum images, we used the same elliptical mask as for the self-calibration procedure. Following the same cleaning strategy used in MAPS (see Czekala et al., 2021, for a detailed description), we defined the robust parameter jointly with a UV taper of the visibilities in order to obtain a common circularized beam of for each transition (see Table 4). The images were CLEANed up to a threshold defined as four times the estimated rms of unmasked regions, and subsequently, were corrected for the “JvM effect” (Jorsater & van Moorsel, 1995). The JvM correction is applied by the parameter, defined as the ratio between the CLEAN and the dirty beam volumes (see Czekala et al., 2021, for more details), whose main effect is a correct accounting of the extended source flux, but which also suppresses the point-source noise in the resulting images. Our derived values are listed in Table 4, where the closer is to , the smaller its effect.
The observational parameters of the final JvM-corrected images are summarized in Table 2. All the final cubes have a channel width of km s, with the exception of the Band 3 transition whose native resolution is about km s. The same procedures where applied then to the continuum measurement sets to obtain the dust continuum images. According to the ALMA Technical Handbook444https://almascience.eso.org/documents-and-tools/cycle10/alma-technical-handbook, the reported flux measurements should include an absolute flux calibration uncertainty of approximately .
@@ 3.1 Disk-integrated Emission
Figure 1 shows some representative emission channels of the brightest HCO transitions and the Keplerian mask used in this work, where we can clearly identify a disk velocity structure consistent with Keplerian rotation. Using the integrated_spectrum function from the Python package GoFish (Teague, 2019), we extracted the disk-integrated spectrum within the Keplerian mask for each transition to obtain their corresponding integrated flux densities, which are listed in Table 2. Following the same bootstrapping procedure described in detail by Bergner et al. (2018) and Pegues et al. (2020), we defined the uncertainty associated to the integrated flux density as the standard deviation of the integrated fluxes measured within 1000 off-source Keplerian masks.
3.2 Spatial Distribution of HCO Emission
To analyze the spatial distribution of HCO in HD 163296, we collapse the spectral cubes into integrated intensity, or “zeroth-moment,” maps using the Python package bettermoments555https://github.com/richteague/bettermoments. The line emission was integrated in velocity using the same Keplerian mask for all transitions (see Figure 1). We did not apply a threshold clipping to recover accurate fluxes. As the number of selected pixels within a Keplerian mask varies with the different velocity channels, then the associated rms is not uniform across the zeroth-moment map (Bergner et al., 2018; Pegues et al., 2020). Thus, we estimated the zeroth-moment rms also using a bootstrapping procedure. We created 1000 zeroth-moment maps, using the same Keplerian mask but centered at different off-source positions. Then, to compute the rms of each pixel, we took the standard deviation of these off-source zeroth-moment maps. Finally, the median value of these standard deviations is considered as the zeroth-moment rms reported in Table 2.

### v2

# Radial and vertical constraints on the icy origin of H2CO in the HD 163296 Protoplanetary Disk

@@ Abstract
H2CO is a small organic molecule widely detected in protoplanetary disks. As a precursor to grain-surface formation of CH3OH, H2CO is considered an important precursor of O-bearing organic molecules that are locked in ices. Still, since gas-phase reactions can also form H2CO, there remains an open question on the channels by which organics form in disks, and how much the grain versus the gas pathways impact the overall organic reservoir. We present spectrally and spatially resolved Atacama Large Millimeter/submillimeter Array observations of several ortho- and para-H2CO transitions toward the bright protoplanetary disk around the Herbig Ae star HD 163296. We derive column density, excitation temperature, and ortho-to-para ratio (OPR) radial profiles for H2CO, as well as disk-averaged values of cm-2, K, and , respectively. We empirically determine the vertical structure of the emission, finding vertical heights of . From the profiles, we find a relatively constant with radius, but still consistent with among the uncertainties, a secondary increase of in the outer disk, and low values that decrease with disk radius. Our resulting radial, vertical, and OPR constraints suggest an increased UV penetration beyond the dust millimeter edge, consistent with an icy origin but also with cold gas-phase chemistry. This Herbig disk contrasts previous results for the T Tauri disk, TW Hya, which had a larger contribution from cold gas-phase chemistry. More observations of other sources are needed to disentangle the dominant formation pathway of H2CO in protoplanetary disks.
@@ 1 Introduction
Although complex hydrocarbons and N-bearing COMs are commonly detected in protoplanetary disks, O-bearing COMs detections are scarce (see Öberg et al., 2023, and references therein) because most of them are expected to be locked in the ice mantles on the surface of dust grains (Öberg et al., 2011; Boogert et al., 2015). Even methanol (CH3OH), one of the simplest O-bearing COMs and considered a building block in forming more complex species (Öberg et al., 2009), has had relatively few detections in disks. Except for TW Hya (Walsh et al., 2016), the closest protoplanetary disk, CH3OH detections are typically restricted to peculiar sources with specific warm conditions (e.g., van ’t Hoff et al., 2018; Lee et al., 2019; Podio et al., 2020) or cavities directly exposed to the radiation of the host star (e.g., Booth et al., 2021; van der Marel et al., 2021; Booth et al., 2023), consistent with an icy origin. Thus, the need arises for an abundant gas-phase molecule capable of tracing the O-rich organic content in protoplanetary disks.
@@
According to previous work (Walsh et al., 2014; Öberg et al., 2017), the small organic molecule formaldehyde (H2CO) has the potential to be a tracer of the frozen content in dust grains since it can efficiently form on ices by the successive hydrogenation of CO, which also leads to the formation of CH3OH and other COMs (e.g., Watanabe & Kouchi, 2002; Chuang et al., 2017) and is more accessible to detect in the gas-phase due to lower desorption energies compared to CH3OH (Carney et al., 2019). Still, H2CO can also be formed through gas-phase reactions (e.g., Fockenberg & Preses, 2002; Atkinson et al., 2006), and thus, it cannot be used as a direct tracer of dust-grain chemical products if there are no constraints on the contribution from each formation pathway (Loomis et al., 2015; Terwisscha van Scheltinga et al., 2021).
@@
Consequently, H2CO has been extensively studied in multiple protoplanetary disks, mostly around low-mass T Tauri stars. As different regions of the disk are expected to have distinct types of chemistry (Öberg & Bergin, 2021), most studies have focused on resolving the radial distribution of H2CO (e.g., Loomis et al., 2015; Öberg et al., 2017; Kastner et al., 2018; Podio et al., 2019; Garufi et al., 2020; Guzmán et al., 2021; Booth et al., 2023). Nevertheless, the disk vertical structure also has an important role in H2CO chemistry, which can be constrained by observing edge-on sources (e.g., Podio et al., 2020), using geometrical methods (e.g., Paneque-Carreño et al., 2023), or determining the excitation conditions by observing multiple H2CO lines (e.g., Pegues et al., 2020).
@@
Parameter Value Reference Distance pc Gaia Collaboration et al. (2018) InclinationaaThe geometric parameters of the disk are derived from the dust millimeter continuum. deg Huang et al. (2018) P.A.aaThe geometric parameters of the disk are derived from the dust millimeter continuum. deg Huang et al. (2018) Teague et al. (2021) Fairlamb et al. (2015) Stellar type A1 Fairlamb et al. (2015) km s-1 Teague et al. (2019)
@@
Moreover, H2CO has two isomeric forms depending on the hydrogen nuclei spin alignment: ortho-H2CO (parallel alignment) and para-H2CO (antiparallel alignment). The ratio of the two isomers, known as the ortho-to-para ratio (OPR), has also been considered a possible way to discriminate between H2CO formation pathways (e.g., Guzmán et al., 2011). However, the interpretation of derived OPRs is still a matter of discussion (Guzmán et al., 2018; Terwisscha van Scheltinga et al., 2021, see Section 4.2.3 for further details). Currently, the excitation conditions and the OPR of H2CO have been well-resolved only for the T Tauri disk TW Hya (Terwisscha van Scheltinga et al., 2021), thanks to the Atacama Large Millimeter/submillimeter Array (ALMA) high angular resolution observations of several ortho and para lines of H2CO. Therefore, further multiple-line, spatially resolved studies of p-H2CO and o-H2CO are needed to gain a representative understanding of H2CO chemistry in protoplanetary disks.
@@
Due to its large size ( au, Law et al., 2021a), proximity, and high luminosity (see Table 1), the Herbig Ae disk HD 163296 is an ideal source for a detailed study of H2CO chemistry. Currently, only two p-H2CO transitions have been spatially resolved in HD 163296 (Carney et al., 2017; Pegues et al., 2020; Guzmán et al., 2021). Still, additional bright p-H2CO and o-H2CO lines can be potentially resolved (Guzmán et al., 2018), thus covering a broad range of excitation energies to obtain well-constrained estimates of the excitation conditions and the OPR of H2CO. The disk also contains a well-defined CO snowline (Qi et al., 2015) and dust millimeter edge (Law et al., 2021a), which makes HD 163296 a perfect target to see how H2CO chemistry works in more massive Herbig Ae/Be sources and compare it to the chemistry in the colder T Tauri, TW Hya disk.
@@
In this work, we analyze the emission of multiple o-H2CO and p-H2CO transitions in HD 163296 resolved at angular resolution with ALMA. We determine empirically the radial and vertical distribution of H2CO, as well as its excitation conditions and OPR as a function of radius. The observations and data reduction are described in Section 2. The results on the spatial distribution, excitation conditions, and OPR are presented in Section 3, while their implications in the formation of H2CO in protoplanetary disks are discussed in Section 4. The conclusions are summarized in Section 5.
@@
H2CO Line Baselines Chan. Width BeamaaThe robust parameter used for Briggs weighting and a UV taper, both applied in the CLEAN process, were computed independently for each transition to obtain a common circularized beam of for all lines (see Appendix A). Beam PA Chan. rmsbbThe channel rms is calculated using the central of the pixels in the first and last five channels of the data cubes. Mom. Zero rmsccThe moment zero rms is defined by the bootstrapping procedure described in Section 3.2. Int. Flux DensityddThe integrated flux density is measured within an outer radius of and integrating the emission between and km s-1. The uncertainties do not include the calibration uncertainty. (m) (km s-1) (′′) (∘) (mJy beam-1) (mJy beam-1 km s-1) (mJy km s-1) ALMA Band 3 o-H2CO 11footnotemark: ALMA Band 6 p-H2CO 11footnotemark: o-H2CO 22footnotemark: ALMA Band 7 p-H2CO 22footnotemark: p-H2CO 22footnotemark: p-H2CO 22footnotemark: o-H2CO 22footnotemark:
@@ 2.1 Observational Details
We present ALMA observations of several H2CO lines in Bands 3, 6, and 7 in the protoplanetary disk HD 163296111HD 163296 protoplanetary disk coordinates are ; .. A detailed description of the Band 6 and 7 observations (project 2016.1.00884.S, PI: V.V. Guzmán) is given in Carney et al. (2019). In summary, the Band 6 observations were carried out with the 12 m array in two different configurations between 2016 November and 2017 March, whereas in the case of Band 7, the short baselines executions were observed with the Atacama Compact Array on 2016 October and the long baselines executions were done with the 12m array on 2017 April. Two different spectral setups were used: one targeting the o-H2CO rotational transition centered at GHz (Band 6), and other targeting several p-H2CO and o-H2CO rotational transitions (, , , and ) centered at GHz (Band 7). Transitions of other molecules were also covered within the above setups, such as CH3OH (Carney et al., 2019) and CN (Bergner et al., 2021).
Additionally, these observations were complemented with publicly available data from the MAPS ALMA Large Program (project 2018.1.01055.L, PI: K.I. Öberg). All the observational details of these data can be found in Öberg et al. (2021). In particular, we used the detected p-H2CO rotational transition at GHz (Band 6) presented by Guzmán et al. (2021), and the nondetected o-H2CO rotational line at GHz (Band 3). All the transitions used in this work are summarized in Table 2 along with their respective project codes.
@@ 2.2 Data Reduction
The data reduction and posterior imaging were performed using the Common Astronomy Software Application (CASA)222https://casa.nrao.edu/ version 6.5.2.26 (McMullin et al., 2007; CASA Team et al., 2022). The data from 2016.1.00884.S were initially processed through the standard pipeline provided by the Joint ALMA Observatory (JAO), and then self-calibrated to improve the signal-to-noise ratio following a methodology similar to that used in the DSHARP ALMA Large Program (see Andrews et al., 2018, for more details). Essentially, we created pseudo-continuum visibilities by flagging the line emission in the observed spectral windows, and then we aligned the different measurement sets (short and long baselines separately) into a common phase center by using the CASA functions fixvis and fixplanets. Afterward, the self-calibration procedure can be summarized as follows: we first applied the phase-only self-calibration to the short baseline visibilities, followed by a phase-amplitude self-calibration. Finally, we repeated the same procedures for the combined measurement sets (short + long baselines).
@@
The calibrated visibilities were imaged through the CLEAN algorithm (Högbom, 1974) implemented by the tclean task. We used a Briggs weighting and multiscale deconvolver (Cornwell, 2008) with scales of 0, 5, 15, and 25 pixels, where the pixel size corresponds to approximately one-seventh of the beam size. The CLEAN mask was defined using the keplerian_mask package (Teague, 2020) to account for the Keplerian rotation of the disk, using the same stellar and disk parameters of HD 163296 as in MAPS (see Öberg et al., 2021, values are summarized in Table 1), ensuring that we covered all disk emission at different velocities. Regarding continuum images, we used the same elliptical mask as for the self-calibration procedure. Following the same cleaning strategy used in MAPS (see Czekala et al., 2021, for a detailed description), we defined the robust parameter jointly with a UV taper of the visibilities in order to obtain a common circularized beam of for each transition (see Table 4). The images were CLEANed up to a threshold defined as four times the estimated rms of unmasked regions, and subsequently, were corrected for the JvM effect (Jorsater & van Moorsel, 1995). The JvM correction is applied by the parameter, defined as the ratio between the CLEAN and the dirty beam volumes (see Czekala et al., 2021, for more details), whose main effect is a correct accounting of the extended source flux, but which also suppresses the point-source noise in the resulting images. Our derived values are listed in Table 4, where the closer is to , the smaller its effect.
The observational parameters of the final JvM-corrected images are summarized in Table 2. All the final cubes have a channel width of km s-1, with the exception of the Band 3 transition, whose native resolution is about km s-1. The same procedures where applied then to the continuum measurement sets to obtain the dust continuum images. According to the ALMA Technical Handbook444https://almascience.eso.org/documents-and-tools/cycle10/alma-technical-handbook, the reported flux measurements should include an absolute flux calibration uncertainty of approximately .
@@ 3.1 Disk-integrated Emission
Figure 1 shows some representative emission channels of the brightest H2CO transitions and the Keplerian mask used in this work, where we can clearly identify a disk velocity structure consistent with Keplerian rotation. Using the integrated_spectrum function from the Python package GoFish (Teague, 2019), we extracted the disk-integrated spectrum within the Keplerian mask for each transition to obtain their corresponding integrated flux densities, which are listed in Table 2. Following the same bootstrapping procedure described in detail by Bergner et al. (2018) and Pegues et al. (2020), we defined the uncertainty associated with the integrated flux density as the standard deviation of the integrated fluxes measured within 1000 off-source Keplerian masks.
3.2 Spatial Distribution of H2CO Emission
To analyze the spatial distribution of H2CO in HD 163296, we collapse the spectral cubes into integrated intensity, or zeroth-moment, maps using the Python package bettermoments555https://github.com/richteague/bettermoments. The line emission was integrated into velocity using the same Keplerian mask for all transitions (see Figure 1). We did not apply a threshold clipping to recover accurate fluxes. As the number of selected pixels within a Keplerian mask varies with the different velocity channels, then the associated rms is not uniform across the zeroth-moment map (Bergner et al., 2018; Pegues et al., 2020). Thus, we estimated the zeroth-moment rms also using a bootstrapping procedure. We created 1000 zeroth-moment maps, using the same Keplerian mask but centered at different off-source positions. Then, to compute the rms of each pixel, we took the standard deviation of these off-source zeroth-moment maps. Finally, the median value of these standard deviations is considered as the zeroth-moment rms reported in Table 2.